# Counting rational points on Hirzebruch–Kleinschmidt varieties over global function fields

### v1

@@ 5.1. Height zeta functions
  1. (1)
    The Riemann–Roch theorem (see Section 2.2) implies that
    $\mathbf{R}_{K}(a,b)=\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\leq 2g-2\end{subarray}}q^{-(a\ell(D)+b\deg(D))}+q^{a(g-1)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)>2g-2\end{subarray}}q^{-(a+b)\deg(D)},$ (14)
    hence $\mathbf{R}_{K}(a,b)=q^{a(g-1)}\zeta_{K}(a+b)+\mathbf{S}_{K}(a,b)$ with the finite sum
    In particular, $\mathbf{R}_{K}(a,b)=q^{-a}\zeta_{K}(a+b)$ if $g=0$.
  2. (2)
    As in the case of Hirzebruch–Kleinschmidt varieties over number fields (see [HMM24, Remark 6.8]), the different cases that appear in Theorem 5.2 give a subdivision of the big cone of $X$, i.e. the interior of $\Lambda_{\textup{eff}}(X)$ (see Figure 1 for an illustration, where we assume $t>|\mathbf{a}|$ for simplicity). It follows from the proof of Theorem 5.2 in Section 5.3 that the line bundles $L$ contained in the ray passing through the anticanonical class have height zeta functions with a double pole at $s=A_{L}=B_{L}$, while line bundles outside this ray have height zeta functions with a simple pole at $s=\max\{A_{L},B_{L}\}$.
$\xi$$\gamma$$-K_{X}$$\xi=\frac{t-|\mathbf{a}|}{r+1}\gamma$$\xi<\frac{t-|\mathbf{a}|}{r+1}\gamma$$\xi>\frac{t-|\mathbf{a}|}{r+1}\gamma$
Figure 1. Subdivision of the big cone of $X_{d}(a_{1},\ldots,a_{r})$ (when $t>|\mathbf{a}|$ for simplicity).
@@
Let $X\simeq\mathbb{P}^{r}\times\mathbb{P}^{t-1}$ be a Hirzebruch–Kleinschmidt variety over the global function field $K=\mathbb{F}_{q}(\mathscr{C})$ with $a_{r}=0$ and let $L=\gamma h+\xi f\in\operatorname{Pic}(X)$ big. Then, the height zeta function $\zeta_{X,L}(s)$ is a rational function in $q^{-s}$. Moreover, $\zeta_{X,L}(s)$ converges absolutely for $\Re(s)>a(L)$ and it has a pole of order $b(L)$ at $s=a(L)$ with
$\lim_{s\to a(L)}(s-a(L))^{b(L)}\zeta_{X,L}(s)=\left\{\begin{array}[]{ll}\frac{q^{(d+2)(1-g)}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)\gamma\xi(q-1)^{2}\log(q)^{2}}&\text{if }\xi=\left(\frac{t}{r+1}\right)\gamma,\vspace{0.1cm}\\ \frac{\zeta_{\mathbb{P}^{r}}(\gamma B_{L})h_{K}q^{t(1-g)}}{\zeta_{K}(t)\xi(q-1)\log(q)}&\text{if }\xi<\left(\frac{t}{r+1}\right)\gamma,\vspace{0.1cm}\\ \frac{\zeta_{\mathbb{P}^{t-1}}(\xi A_{L})h_{K}q^{(r+1)(1-g)}}{\gamma\zeta_{K}(r+1)(q-1)\log(q)}&\text{if }\xi>\left(\frac{t}{r+1}\right)\gamma.\vspace{0.1cm}\end{array}\right.$
@@
    If $A_{L}<B_{L}$, then $\zeta_{X,L}(s)$ has a simple pole at $s=B_{L}=a(L)$ with
    $\begin{split}\lim_{s\to a(L)}(s-a(L))\zeta_{X,L}(s)&=\frac{\zeta_{\mathbb{P}^{r}}(\gamma B_{L})\operatorname{Res}_{s=t}\zeta_{\mathbb{P}^{t-1}}(s)}{\xi}\\ &=\frac{\zeta_{\mathbb{P}^{r}}(\gamma B_{L})h_{K}q^{t(1-g)}}{\xi\zeta_{K}(t)(q-1)\log(q)}.\end{split}$
@@
    If $A_{L}>B_{L}$, then $\zeta_{X,L}(s)$ has a simple pole at $s=A_{L}=a(L)$ with
    $\begin{split}\lim_{s\to a(L)}(s-a(L))\zeta_{X,L}(s)&=\frac{\operatorname{Res}_{s=r+1}\zeta_{\mathbb{P}^{r}}(s)\zeta_{\mathbb{P}^{t-1}}(\xi A_{L})}{\gamma}\\ &=\frac{\zeta_{\mathbb{P}^{t-1}}(\xi A_{L})h_{K}q^{(r+1)(1-g)}}{\gamma\zeta_{K}(r+1)(q-1)\log(q)}.\end{split}$
These cases correspond exactly to $\xi$ equal, less than, or greater than $\left(\frac{t}{r+1}\right)\gamma$, respectively. This proves the theorem. ∎
@@ 5.2. Preliminary lemmas
In this section we stablish some preliminary lemmas needed for the proof of Theorem 5.2 in the next section.
@@
Moreover, we define the functions $1,\delta,\mu:\operatorname{Div}^{+}(K)\rightarrow\mathbb{C}$ by
@@
  • $\delta(D):=\left\{\begin{array}[]{ll}1&\textup{if }D=0,\\ 0&\textup{otherwise}.\end{array}\right.$
  • $\mu(D):=\left\{\begin{array}[]{ll}1&\textup{if }D=0,\\ (-1)^{\sum_{v}v(D)}&\textup{if $D\neq 0$ and $v(D)=0$ or $1$ for all $v\in\operatorname{Val}(K)$},\\ 0&\textup{otherwise}.\end{array}\right.$
In particular, $\delta$ is a unit for the convolution product and $1\star\mu=\mu\star 1=\delta$.
@@
Let $D\in\operatorname{Div}^{+}(K)$ and let $f,g:\operatorname{Div}^{+}(K)\to\mathbb{C}$ be two functions. Then, the relation
$f(D)=\sum_{0\leq D^{\prime}\leq D}g(D^{\prime})$
@@
$g(D)=\sum_{0\leq D^{\prime}\leq D}\mu(D-D^{\prime})f(D^{\prime}).$
@@
By definition of the function $\delta$, the pair $(\delta,1)$ forms a $\mu$-couple.
@@
Since $(\delta,1)$ is a $\mu$-couple, we have that
$\begin{split}\sum_{D^{\prime}\geq 0}\mu(D^{\prime})T^{\deg(D^{\prime})}\operatorname{Z}_{K}(T)&=\sum_{D^{\prime}\geq 0}\mu(D^{\prime})T^{\deg(D^{\prime})}\sum_{D\geq 0}T^{\deg(D)}\\ &=\sum_{D\geq 0}\sum_{D^{\prime}\geq 0}\mu(D^{\prime})T^{\deg(D+D^{\prime})}\\ &=\sum_{D\geq 0}\sum_{0\leq D^{\prime}\leq D}\mu(D-D^{\prime})T^{\deg(D)}\\ &=\sum_{D\geq 0}\delta(D)T^{\deg(D)}=1,\end{split}$
where the last equality follows from the definition of the function $\delta$. This proves the lemma. ∎
@@
The good open subset $U_{d}(a_{1},\ldots,a_{r})$ equals the image of the map
@@
With the above notation, we have the equality
@@
The following relation holds:
@@
$N_{m}^{n}(D):=\#\left\{(x_{i})\in\mathbb{A}^{m}:n\sup_{i\in I_{m}}(x_{i})_{\infty}\leq D\right\},$
@@
$\tilde{N}_{m}(D):=\#\left\{(x_{i})\in\mathbb{A}^{m}:\sup_{i\in I_{m}}(x_{i})_{\infty}=D\right\}.$
@@
with $A_{L},B_{L}$ defined in (12). Finally, recalling that $r\geq 1,t\geq 2$, we get the following properties:
@@
    $\begin{split}\lim_{s\to a(L)}(s-a(L))^{2}\zeta_{U,L}(s)&=\frac{q^{d(1-g)}}{\zeta_{K}(t)\zeta_{K}(r+1)}\frac{\big{(}\textup{Res}_{s=1}\zeta_{K}(s)\big{)}^{2}}{(\gamma+a_{r}\xi)\gamma}\\ &=\frac{q^{(d+2)(1-g)}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)(\gamma+a_{r}\xi)\gamma(q-1)^{2}\log(q)^{2}}.\end{split}$
@@
    If $A_{L}<B_{L}$, then $\zeta_{U,L}(s)$ satisfies
    $\begin{split}\lim_{s\to a(L)}(s-a(L))\zeta_{U,L}(s)&=\left(q^{d(1-g)}\zeta_{K}(\gamma B_{L}-r)+q^{(d+1-N_{X})(1-g)}P_{2}(q^{-B_{L}})\right)\\ &\hskip 56.9055pt\times\frac{\textup{Res}_{s=1}\zeta_{K}(s)}{\zeta_{K}\left(t\right)\zeta_{K}(\gamma B_{L})(\gamma a_{r}+\xi)}\\ &=\left(q^{d(1-g)}\zeta_{K}(\gamma B_{L}-r)+q^{(d+1-N_{X})(1-g)}P_{2}(q^{-B_{L}})\right)\\ &\hskip 56.9055pt\times\frac{h_{K}q^{1-g}}{\zeta_{K}\left(t\right)\zeta_{K}(\gamma B_{L})(\gamma a_{r}+\xi)(q-1)\log(q)}.\end{split}$
    From the definition of $P_{2}(T)$ and (14) we have
    $\begin{split}&q^{d(1-g)}\zeta_{K}(\gamma B_{L}-r)+q^{(d+1-N_{X})(1-g)}P_{2}(q^{-B_{L}})\\ =&q^{(d+1-N_{X})(1-g)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\leq 2g-2\end{subarray}}q^{(N_{X}-1)\ell(D)+(r+1-N_{X}-\gamma B_{L})\deg(D)}+q^{d(1-g)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)>2g-2\end{subarray}}q^{-(\gamma B_{L}-r)\deg(D)}\\ =&q^{(d+1-N_{X})(1-g)}\mathbf{R}_{K}(1-N_{X},\gamma B_{L}-r+N_{X}-1).\end{split}$
    Since this value is positive, we conclude that $\zeta_{U,L}(s)$ has a simple pole at $s=a(L)$ in this case.
@@
    If $A_{L}>B_{L}$, then $\zeta_{U,L}(s)$ satisfies
    $\begin{split}\lim_{s\to a(L)}(s-a(L))\zeta_{U,L}(s)&=\left(q^{d(1-g)}\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)A_{L}-\left((r+1)a_{r}-|\mathbf{a}|+t-1\right)\right)+q^{r(1-g)}P_{3}(q^{-A_{L}})\right)\\ &\hskip 56.9055pt\times\frac{\textup{Res}_{s=1}\zeta_{K}(s)}{\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)A_{L}-\left((r+1)a_{r}-|\mathbf{a}|\right)\right)\zeta_{K}(r+1)\gamma}\\ &=\left(q^{d(1-g)}\zeta_{K}\left(A_{L}\xi+|\mathbf{a}|-t+1\right)+q^{r(1-g)}P_{3}(q^{-A_{L}})\right)\\ &\hskip 56.9055pt\times\frac{h_{K}q^{1-g}}{\zeta_{K}\left(A_{L}\xi+|\mathbf{a}|\right)\zeta_{K}(r+1)\gamma(q-1)\log(q)}.\end{split}$
@@
    $\begin{split}&q^{d(1-g)}\zeta_{K}\left(A_{L}\xi+|\mathbf{a}|-t+1\right)+q^{r(1-g)}P_{3}(q^{-A_{L}})\\ =&q^{r(1-g)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\leq 2g-2\end{subarray}}q^{\ell(D)(t-1)-\deg(D)\left(A_{L}\xi+|\mathbf{a}|\right)}+q^{d(1-g)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)>2g-2\end{subarray}}q^{-\left(A_{L}\xi+|\mathbf{a}|-t+1\right)\deg(D)}\\ =&q^{r(1-g)}\mathbf{R}_{K}\left(1-t,A_{L}\xi+|\mathbf{a}|\right).\end{split}$

### v2

@@ 5.1. Height zeta functions
As in the case of Hirzebruch–Kleinschmidt varieties over number fields (see [HMM24, Remark 6.8]), the different cases that appear in Theorem 5.2 give a subdivision of the big cone of $X$, i.e. the interior of $\Lambda_{\textup{eff}}(X)$ (see Figure 1 for an illustration, where we assume $t>|\mathbf{a}|$ for simplicity). Indeed, the line bundles $L$ contained in the ray passing through the anticanonical class have height zeta functions with a double pole at $s=A_{L}=B_{L}$, while line bundles outside this ray have height zeta functions with a simple pole at $s=\max\{A_{L},B_{L}\}$.
[…]
Figure 1. Subdivision of the big cone of $X_{d}(a_{1},\ldots,a_{r})$ according to the values of $A_{L}=\frac{r+1}{\gamma}$ and $B_{L}=\frac{(r+1)a_{r}-|\mathbf{a}|+t}{\gamma a_{r}+\xi}$ (assuming $t>|\mathbf{a}|$ for simplicity).
@@
Let $X\simeq\mathbb{P}^{r}\times\mathbb{P}^{t-1}$ be a Hirzebruch–Kleinschmidt variety over the global function field $K=\mathbb{F}_{q}(\mathscr{C})$ with $a_{r}=0$ and let $L=\gamma h+\xi f\in\operatorname{Pic}(X)$ big and primitive. Then, $\zeta_{X,L}(s)$ converges absolutely for $\Re(s)>a(L)$ and is a rational function in $q^{-s}$. Moreover, the following properties hold:
  1. (1)
    If $A_{L}=B_{L}$, then $\zeta_{X,L}(s)$ is holomorphic for
    $\Re(s)>\max\left\{\frac{1}{2\gamma},\frac{1}{2\xi}\right\},s\not\in\left\{w+\frac{2\pi im}{\log(q)}:w\in\{a(L)\}\cup\mathscr{A}_{L}\cup\mathscr{B}_{L},m\in\mathbb{Z}\right\},$
    it has a pole of order $b(L)=2$ at $s=a(L)$ and satisfies
    $\lim_{s\to w}(s-w)^{2}\zeta_{X,L}(s)=\left\{\begin{array}[]{ll}\frac{q^{(d+2)(1-g)}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)\xi\gamma(q-1)^{2}\log(q)^{2}}&\text{if }w=a(L),\\ 0&\text{if }w\in\mathscr{A}_{L}\cup\mathscr{B}_{L}.\end{array}\right.$
  2. (2)
    If $A_{L}<B_{L}$, then $\zeta_{X,L}(s)$ is holomorphic for
    $\Re(s)>\max\left\{A_{L},\frac{1}{2\xi}\right\},s\not\in\left\{w+\frac{2\pi im}{\log(q)}:w\in\{a(L)\}\cup\mathscr{B}_{L},m\in\mathbb{Z}\right\},$
    it has a pole of order $b(L)=1$ at $s=a(L)$ and for every $w\in\{a(L)\}\cup\mathscr{B}_{L}$ it satisfies
    $\lim_{s\to w}(s-w)\zeta_{X,L}(s)=\frac{\zeta_{\mathbb{P}^{r}}(\gamma w)h_{K}q^{t(1-g)}}{\zeta_{K}(t)\xi(q-1)\log(q)}.$
  3. (3)
    If $A_{L}>B_{L}$, then $\zeta_{X,L}(s)$ is holomorphic for
    $\Re(s)>\max\left\{\frac{1}{2\gamma},B_{L}\right\},s\not\in\left\{w+\frac{2\pi im}{\log(q)}:w\in\{a(L)\}\cup\mathscr{A}_{L},m\in\mathbb{Z}\right\},$
    it has a pole of order $b(L)=1$ at $s=a(L)$ and for every $w\in\{a(L)\}\cup\mathscr{A}_{L}$ it satisfies
    $\lim_{s\to w}(s-w)\zeta_{X,L}(s)=\frac{\zeta_{\mathbb{P}^{t-1}}(\xi w)h_{K}q^{(r+1)(1-g)}}{\gamma\zeta_{K}(r+1)(q-1)\log(q)}.$
@@
    Moreover, $\zeta_{U,L}(s)$ is holomorphic for
    $\Re(s)>\max\left\{\frac{1}{2\gamma},\frac{1}{2\xi}\right\},s\not\in\left\{w+\frac{2\pi im}{\log(q)}:w\in\{a(L)\}\cup\mathscr{A}_{L}\cup\mathscr{B}_{L},m\in\mathbb{Z}\right\},$
    and satisfies $\lim_{s\to w}(s-w)^{2}\zeta_{U,L}(s)=0$ for every $w\in\mathscr{A}_{L}\cup\mathscr{B}_{L}$, since $\mathscr{A}_{L}\cap\mathscr{B}_{L}=\emptyset$.
@@
    If $A_{L}<B_{L}$, then $\zeta_{X,L}(s)$ is holomorphic for
    $\Re(s)>\max\left\{A_{L},\frac{1}{2\xi}\right\},s\not\in\left\{w+\frac{2\pi im}{\log(q)}:w\in\{a(L)\}\cup\mathscr{B}_{L},m\in\mathbb{Z}\right\},$
    and for every $w\in\{a(L)\}\cup\mathscr{B}_{L}$ it satisfies
    $\begin{split}\lim_{s\to w}(s-w)\zeta_{X,L}(s)&=\frac{\zeta_{\mathbb{P}^{r}}(\gamma w)\operatorname{Res}_{s=t}\zeta_{\mathbb{P}^{t-1}}(s)}{\xi}\\ &=\frac{\zeta_{\mathbb{P}^{r}}(\gamma w)h_{K}q^{t(1-g)}}{\xi\zeta_{K}(t)(q-1)\log(q)}.\end{split}$
    In particular, since $\zeta_{\mathbb{P}^{r}}(\gamma B_{L})$ is positive, $\zeta_{X,L}(s)$ has a pole of order $b(L)=1$ at $s=a(L)$.
@@
    If $A_{L}>B_{L}$, then $\zeta_{X,L}(s)$ is holomorphic for
    $\Re(s)>\max\left\{\frac{1}{2\gamma},B_{L}\right\},s\not\in\left\{w+\frac{2\pi im}{\log(q)}:w\in\{a(L)\}\cup\mathscr{A}_{L},m\in\mathbb{Z}\right\},$
    and for every $w\in\{a(L)\}\cup\mathscr{A}_{L}$ it satisfies
    $\begin{split}\lim_{s\to w}(s-w)\zeta_{X,L}(s)&=\frac{\operatorname{Res}_{s=r+1}\zeta_{\mathbb{P}^{r}}(s)\zeta_{\mathbb{P}^{t-1}}(\xi w)}{\gamma}\\ &=\frac{\zeta_{\mathbb{P}^{t-1}}(\xi w)h_{K}q^{(r+1)(1-g)}}{\gamma\zeta_{K}(r+1)(q-1)\log(q)}.\end{split}$
    As in the previous case, using that $\zeta_{\mathbb{P}^{t-1}}(\xi A_{L})$ is positive, we conclude that $\zeta_{X,L}(s)$ has a pole of order $b(L)=1$ at $s=a(L)$.
This proves the theorem. ∎
@@ 5.2. Preliminary lemmas
In this section we establish some preliminary lemmas needed for the proof of Theorem 5.2 in the next section.
@@
Moreover, we define the functions $1,u,\mu:\operatorname{Div}^{+}(K)\rightarrow\mathbb{C}$ by
@@
  • $u(D):=\left\{\begin{array}[]{ll}1&\textup{if }D=0,\\ 0&\textup{otherwise}.\end{array}\right.$
  • $\mu(D):=\left\{\begin{array}[]{ll}(-1)^{\sum_{v}v(D)}&\textup{if~{}$v(D)=0$ or$1$ for all $v\in\operatorname{Val}(K)$},\\ 0&\textup{otherwise}.\end{array}\right.$
In particular, $u$ is a unit for the convolution product and $1\star\mu=\mu\star 1=u$.
@@
Let $f,g:\operatorname{Div}^{+}(K)\to\mathbb{C}$ be two functions. Then, the relation
$f(D)=\sum_{0\leq D^{\prime}\leq D}g(D^{\prime})\quad\text{for all }D\in\operatorname{Div}^{+}(K)$
@@
$g(D)=\sum_{0\leq D^{\prime}\leq D}\mu(D-D^{\prime})f(D^{\prime})\quad\text{forall }D\in\operatorname{Div}^{+}(K).$
@@
By definition of the function $u$, the pair $(1,u)$ forms a $\mu$-couple.
@@
Since $(1,u)$ is a $\mu$-couple, we have that
$\begin{split}\sum_{D^{\prime}\geq 0}\mu(D^{\prime})T^{\deg(D^{\prime})}\operatorname{Z}_{K}(T)&=\sum_{D^{\prime}\geq 0}\mu(D^{\prime})T^{\deg(D^{\prime})}\sum_{D\geq 0}T^{\deg(D)}\\ &=\sum_{D\geq 0}\sum_{D^{\prime}\geq 0}\mu(D^{\prime})T^{\deg(D+D^{\prime})}\\ &=\sum_{D\geq 0}\sum_{0\leq D^{\prime}\leq D}\mu(D-D^{\prime})T^{\deg(D)}\\ &=\sum_{D\geq 0}u(D)T^{\deg(D)}=1,\end{split}$
where the last equality follows from the definition of the function $u$. This proves the lemma. ∎
@@
Assume $a_{r}>0$. Then, the good open subset $U_{d}(a_{1},\ldots,a_{r})$ equals the image of the map
@@
Assume $a_{r}>0$. Then, with the above notation, we have the equality
@@
Assuming $a_{r}>0$, the following relation holds:
@@
$N_{m}^{n}(D):=\#\left\{(x_{i})\in\mathbb{A}^{m}:n\sup_{i\in I_{m}}(x_{i})_{\infty}\leq D\right\},$ (23)
@@
$\tilde{N}_{m}(D):=\#\left\{(x_{i})\in\mathbb{A}^{m}:\sup_{i\in I_{m}}(x_{i})_{\infty}=D\right\}.$ (24)
@@
with $A_{L},B_{L}$ defined in (18). Moreover, noting that the sets $\mathscr{A}_{L}$ and $\mathscr{B}_{L}$ are disjoint (because $\mathrm{g.c.d.}(\gamma,\gamma a_{r}+\xi)=\mathrm{g.c.d.}(\gamma,\xi)=1$), we get the following properties:
@@
    $\begin{split}\lim_{s\to a(L)}(s-a(L))^{2}\zeta_{U,L}(s)&=\frac{q^{d(1-g)}}{\zeta_{K}(t)\zeta_{K}(r+1)}\frac{\big{(}\textup{Res}_{s=1}\zeta_{K}(s)\big{)}^{2}}{(\gamma a_{r}+\xi)\gamma}\\ &=\frac{q^{(d+2)(1-g)}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)(\gamma a_{r}+\xi)\gamma(q-1)^{2}\log(q)^{2}}.\end{split}$
    Moreover, $\zeta_{U,L}(s)$ is holomorphic for
    $\Re(s)>\max\left\{A_{L}-\frac{1}{\gamma},B_{L}-\frac{1}{\gamma a_{r}+\xi}\right\},s\not\in\left\{w+\frac{2\pi im}{\log(q)}:w\in\{a(L)\}\cup\mathscr{A}_{L}\cup\mathscr{B}_{L},m\in\mathbb{Z}\right\},$
    and satisfies $\lim_{s\to w}(s-w)^{2}\zeta_{U,L}(s)=0$ for every $w\in\mathscr{A}_{L}\cup\mathscr{B}_{L}$.
@@
    If $A_{L}<B_{L}$, then $\zeta_{U,L}(s)$ is holomorphic for
    $\Re(s)>\max\left\{A_{L},B_{L}-\frac{1}{\gamma a_{r}+\xi}\right\},s\not\in\left\{w+\frac{2\pi im}{\log(q)}:w\in\{a(L)\}\cup\mathscr{B}_{L},m\in\mathbb{Z}\right\},$
    and for every $w\in\{a(L)\}\cup\mathscr{B}_{L}$ it satisfies
    $\begin{split}\lim_{s\to w}(s-w)\zeta_{U,L}(s)&=\left(q^{d(1-g)}\zeta_{K}(\gamma w-r)+q^{(d+1-N_{X})(1-g)}P_{2}(q^{-w})\right)\\ &\hskip 56.9055pt\times\frac{\textup{Res}_{s=1}\zeta_{K}(s)}{\zeta_{K}\left(t\right)\zeta_{K}(\gamma w)(\gamma a_{r}+\xi)}\\ &=\left(q^{d(1-g)}\zeta_{K}(\gamma w-r)+q^{(d+1-N_{X})(1-g)}P_{2}(q^{-w})\right)\\ &\hskip 56.9055pt\times\frac{h_{K}q^{1-g}}{\zeta_{K}\left(t\right)\zeta_{K}(\gamma w)(\gamma a_{r}+\xi)(q-1)\log(q)}.\end{split}$
    Moreover, from the definition of $P_{2}(T)$ and (20) we have
    $\begin{split}&q^{d(1-g)}\zeta_{K}(\gamma w-r)+q^{(d+1-N_{X})(1-g)}P_{2}(q^{-w})\\ =&q^{(d+1-N_{X})(1-g)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\leq 2g-2\end{subarray}}q^{(N_{X}-1)\ell(D)+(r+1-N_{X}-\gamma w)\deg(D)}+q^{d(1-g)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)>2g-2\end{subarray}}q^{-(\gamma w-r)\deg(D)}\\ =&q^{(d+1-N_{X})(1-g)}\mathbf{R}_{K}(1-N_{X},\gamma w-r+N_{X}-1).\end{split}$
    In particular, since $\mathbf{R}_{K}(1-N_{X},\gamma B_{L}-r+N_{X}-1)$ is positive, we see that $\zeta_{U,L}(s)$ has a pole of order $b(L)=1$ at $s=a(L)$.
@@
    If $A_{L}>B_{L}$, then $\zeta_{U,L}(s)$ is holomorphic for
    $\Re(s)>\max\left\{A_{L}-\frac{1}{\gamma},B_{L}\right\},s\not\in\left\{w+\frac{2\pi im}{\log(q)}:w\in\{a(L)\}\cup\mathscr{A}_{L},m\in\mathbb{Z}\right\},$
    and for every $w\in\{a(L)\}\cup\mathscr{A}_{L}$ it satisfies
    $\begin{split}\lim_{s\to w}(s-w)\zeta_{U,L}(s)&=\left(q^{d(1-g)}\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)w-\left((r+1)a_{r}-|\mathbf{a}|+t-1\right)\right)+q^{r(1-g)}P_{3}(q^{-w})\right)\\ &\hskip 56.9055pt\times\frac{\textup{Res}_{s=1}\zeta_{K}(s)}{\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)w-\left((r+1)a_{r}-|\mathbf{a}|\right)\right)\zeta_{K}(r+1)\gamma}\\ &=\left(q^{d(1-g)}\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)w-\left((r+1)a_{r}-|\mathbf{a}|+t-1\right)\right)+q^{r(1-g)}P_{3}(q^{-w})\right)\\ &\hskip 56.9055pt\times\frac{h_{K}q^{1-g}}{\zeta_{K}\left(\left(\gamma a_{r}+\xi\right)w-\left((r+1)a_{r}-|\mathbf{a}|\right)\right)\zeta_{K}(r+1)\gamma(q-1)\log(q)}.\end{split}$
@@
    $\begin{split}&q^{d(1-g)}\zeta_{K}((\gamma a_{r}+\xi)w-((r+1)a_{r}-|\mathbf{a}|+t-1))+q^{r(1-g)}P_{3}(q^{-w})\\ =&q^{r(1-g)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)\leq 2g-2\end{subarray}}q^{\ell(D)(t-1)-\deg(D)((\gamma a_{r}+\xi)w-((r+1)a_{r}-|\mathbf{a}|))}\\ &\hskip 170.71652pt+q^{d(1-g)}\sum_{\begin{subarray}{c}D\geq 0\\ \deg(D)>2g-2\end{subarray}}q^{-((\gamma a_{r}+\xi)w-((r+1)a_{r}-|\mathbf{a}|+t-1))\deg(D)}\\ =&q^{r(1-g)}\mathbf{R}_{K}\left(1-t,\left(\gamma a_{r}+\xi\right)w-((r+1)a_{r}-|\mathbf{a}|)\right).\end{split}$